OpenStudy (anonymous):

\(R\), a closed, bounded, connected region in the plane, moves at a rate of 2 units per second in the direction of \(4\vec{i}+3\vec{j}\). At time \(t\) seconds, with \(-\infty

OpenStudy (anonymous):

\[A(t)=\begin{cases} 0&\text{for }t<=5\\\\ -\dfrac{16}{25}t^2+\dfrac{124}{5}t-108&\text{for }5\le t\le\dfrac{15}{2}\\\\ -\dfrac{16}{5}t+66&\text{for }\dfrac{15}{2}\le t\le15\\\\ \dfrac{16}{25}t^2-28t+294&\text{for }15\le t\le\dfrac{35}{2}\\\\ 0&\text{for }t\ge\dfrac{35}{2} \end{cases}\] Find the area of \(R\).

OpenStudy (jhannybean):

Your alignment skills are impeccable.